# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ III Model calculations
Event generators are the most crucial computational tools used in contemporary high-energy physics phenomenology, enabling to bridge the gap between theoretical models and experimental data. These act as the most effective substitutes for realistic experiments involving hadronic and nucleus-nucleus collisions (AA). These generators are based on the Monte Carlo techniques for simulating the collisions using physics processes to better understand the nature of the events. Based on the kind of colliding species involved and underlying known physics processes, there are several event generators like AMPT, PYTHIA, EPOS LHC, EPOS4, and HERWIG etc, modeled for hadronic and nuclear collisions. In this study, we employ PYTHIA8 and AMPT models as event generators to explore the possible correlation between all event shape classifiers discussed in Sec. II. The detailed mechanism and event generation using these models are described below.
III.1 For $pp$ collisions: PYTHIA 8
PYTHIA is a pQCD-inspired general-purpose Monte Carlo event generator used to simulate relativistic hadronic, leptonic, and heavy-ion collisions from RHIC to LHC energies. In PYTHIA8, models and theories are incorporated for a number of physics aspects, such as parton distributions, hard and soft interactions, initial- and final-state parton showers, multiparton interactions, fragmentation, partonic and hadronic decays [125, 124]. A detailed explanation of all physics processes involved in PYTHIA8 can be found in Ref. [126, 127].
@@ III.1 For $pp$ collisions: PYTHIA 8
In the present study, we use PYTHIA 8.308, an improved version of PYTHIA6, which includes the multi-partonic interactions scenario as one of the important improvements. Using the standard Monash 2013 Tune (Tune:pp = 14) [128], we generate 60 million events in $pp$ collisions at $\sqrt{s}$ = 13 TeV for this analysis. In this analysis, the inelastic and soft QCD events (SoftQCD:inelastic=on) are simulated, as a result, the total scattering cross section includes the contribution from all the single, double, and central diffractive components. We have considered the MPI (PartonLevel:MPI = on) along with mode 1 of color reconnection (ColourReconnection:mode = 1) and Beam Remnants (BeamRemnants:remnantMode = 1). The mode 1 of CR is a newer QCD-based scheme which builds on the dipole formation via the QCD color rules as well as the minimization of the string length. This scheme introduces the junction structure, as a result of which junctions are produced between three or four dipoles. A detailed mechanism can be found in Ref. [127]. Additionally, along with the CR mechanism, we employ another hadronization mechanism for color strings known as rope hadronization (RH). The physical description of CR and RH can be found below.
III.1.1 Color Reconnection
@@
The hadronisation through the fragmentation-based models, such as Lund’s string fragmentation model [129], can be envisaged as a stretched color flux tube between two partons originating from individual partonic scatterings leading to linear confinements through a massless relativistic string. With the increase in the potential energy of the strings, the quark-antiquark pair moves apart till a new quark-antiquark pair is formed. The hadrons can be formed by combining these quarks and antiquarks. The hadronisation through fragmentation in this scenario is independent of each individual scattering, also shown in Fig. 2 (a). However, in the presence of color reconnection, the strings connecting the partonic endpoints from each individual scattering can color reconnect [130]. This is clearly depicted in Fig. 2 (b). Therefore, the fragmentation of two individual scatterings is dependent on each other when the color reconnection is considered, which induces a rise in average transverse momentum with an increase in the number of individual scatterings or the number of multi-partonic interactions.
III.1.2 Rope Hadronisation
@@
The RH mechanism of PYTHIA8 has various tuning parameters. In this study, we use similar settings to the string-shoving mechanism introduced in PYTHIA8. The values of the parameters used for rope hadronization are shown in Table 1. Furthermore, the flag partonvertex (PartonVertex:setVertex = on) is used to set impact-parameter plane vertices for partonic production by ISR, FSR, MPI, and beam remnants. In the simulated events, the hadronic level decay mode (HadronLevel:Decay = on) is enabled for all the resonances except the ones used in our study.
To check the compatibility of PYTHIA8 with experimental data, we have used the same tuning as used in one of our previous works described in Ref [60], where we have compared the transverse momentum and pseudorapidity spectra obtained from PYTHIA8 with the ALICE experimental data for all charged particles in $pp$ collisions at $\sqrt{s}$ = 13 TeV and found them to be comparable within uncertainties.
@@
Table 1: The parameter values of the rope hadronization model used with color reconnection mechanism.
III.2 For Pb–Pb collisions: AMPT
A Multi-Phase Transport (AMPT) is a hybrid model that combines both the initial partonic collisions and the final hadronic interactions effectively, as well as accounts for the transition between these two distinct phases of matter. It has four major components, namely, initial conditions, partonic interactions, conversion of partonic to hadronic matter through the hadronization mechanism, and hadron interactions. In the present study, we use the string melting (SM) mode of AMPT (version 2.26t9b) as the particle anisotropic flow coefficients and spectra are well described at the intermediate-$p_{\rm T}$ region by the quark coalescence mechanism for hadronization [132, 131].
@@ III.2 For Pb–Pb collisions: AMPT
    Initial conditions: The AMPT model employs the heavy-ion jet interaction generator (HIJING)[133] for initial conditions. In the HIJING model, the differential scattering cross section of the energetic minijet partons and soft strings with excited strings, which are further converted to partons, are calculated in $pp$ collisions and then parametrized into the heavy-ion collisions. These are incorporated via the nuclear overlap and shadowing function using an inbuilt Glauber model.
@@
    Parton transport: The produced partons are then transported into the parton transport part called Zhang’s Parton Cascade (ZPC) model [134]. In the String Melting version of AMPT (AMPT-SM), colored strings melt into low-momentum partons.
@@
    Hadronization: A spatial coalescence mechanism [135, 84] is used for the hadronization of the transported partons in AMPT-SM, whereas in the default AMPT version, a Lund string fragmentation mechanism is used for hadronization of the transported partons.
@@
    Hadron transport: The final evolution of produced hadrons is done using a relativistic transport mechanism through meson-meson, meson-baryon, and baryon-baryon interactions [136, 137].

### v2

@@ 9 Appendix
Event generators are the most crucial computational tools used in contemporary high-energy physics phenomenology, bridging the gap between theoretical models and experimental data. These act as the most effective substitutes for realistic experiments involving hadronic and nucleus-nucleus collisions (AA). These generators are based on the Monte Carlo techniques for simulating the collisions using physics processes to better understand the nature of the events. Based on the kind of colliding species involved and underlying known physics processes, there are several event generators like AMPT, PYTHIA, EPOS LHC, EPOS4, and HERWIG, etc, modeled for hadronic and nuclear collisions. We employ PYTHIA8 and AMPT models as event generators to explore the possible correlation between all event shape classifiers discussed in Sec. 2. The detailed mechanism and event generation using these models are described below.
9.1 PYTHIA 8
@@
In the present study, we use PYTHIA 8.308, an improved version of PYTHIA6, which includes the multi-partonic interactions scenario as one of the important improvements. Using the standard Monash 2013 Tune (Tune:pp = 14) Skands:2014pea , we generate 60 million events in $pp$ collisions at $\sqrt{s}$ = 13 TeV for this analysis. The inelastic and soft QCD events (SoftQCD:inelastic=on) are simulated; as a result, the total scattering cross section includes the contribution from all the single, double, and central diffractive components. We have considered the MPI (PartonLevel:MPI = on) along with mode 1 of color reconnection (ColourReconnection:mode = 1) and Beam Remnants (BeamRemnants:remnantMode = 1). The mode 1 of CR is a newer QCD-based scheme which builds on the dipole formation via the QCD color rules as well as the minimization of the string length. This scheme introduces the junction structure, as a result of which junctions are produced between three or four dipoles. A detailed mechanism can be found in Ref. manual . Additionally, along with the CR mechanism, we employ another hadronization mechanism for color strings known as rope hadronization (RH). The physical description of CR and RH can be found below.
9.1.1 Color Reconnection
@@
The hadronisation through the fragmentation-based models, such as Lund’s string fragmentation model Andersson:1983jt , can be envisaged as a stretched color flux tube between two partons originating from individual partonic scatterings leading to linear confinements through a massless relativistic string. With the increase in the potential energy of the strings, the quark-antiquark pair moves apart till a new quark-antiquark pair is formed. The hadrons can be formed by combining these quarks and antiquarks. The hadronisation through fragmentation in this scenario is independent of each individual scattering, also shown in Fig. 2 (a). However, in the presence of color reconnection, the strings connecting the partonic endpoints from each individual scattering can color reconnect Argyropoulos:2014zoa . This is clearly depicted in Fig. 2 (b). Therefore, the fragmentation of two individual scatterings is dependent on each other when the color reconnection is considered, which induces a rise in average transverse momentum with an increase in the number of individual scatterings or the number of multi-partonic interactions.
9.1.2 Rope Hadronisation
@@
The RH mechanism of PYTHIA8 has various tuning parameters. We use similar settings to the string-shoving mechanism introduced in PYTHIA8. The values of the parameters used for rope hadronization are shown in Table 1. Furthermore, the flag partonvertex (PartonVertex:setVertex = on) is used to set impact-parameter plane vertices for partonic production by ISR, FSR, MPI, and beam remnants. In the simulated events, the hadronic level decay mode (HadronLevel:Decay = on) is enabled for all the resonances except the ones used in our study.
To check the compatibility of PYTHIA8 with experimental data, we have used the same tuning as used in one of our previous works described in Ref Prasad:2024gqq , where we have compared the transverse momentum and pseudorapidity spectra obtained from different tunes of PYTHIA8 with the ALICE experimental data ALICE:2015qqj for all charged particles in $pp$ collisions at $\sqrt{s}$ = 13 TeV, as shown in Fig. 86. It can be observed that PYTHIA8 with Color Ropes and Monash show a better quantitative and qualitative agreement to experimental data than PYTHIA8 with Monash NoCR.
@@
Table 1: The parameter values of the rope hadronization model used with the color reconnection mechanism.
Figure 86: Transverse momentum (left) and pseudorapidity (right) differential yield distribution in minimum bias pp collisions at $\sqrt{s}=13$ TeV using PYTHIA8 Color Ropes, Monash and Monash NoCR compared with corresponding measurements with ALICE ALICE:2015qqj .
Apart from the above-mentioned PYTHIA tunes, there are a few additional MC tunes, such as PYTHIA CP1, CP5, A3, A14, etc. are used in CMS and ATLAS experiments. These new tunes differ according to the order of the parton distribution functions (PDF) set used at leading (LO), next-to-leading (NLO), or next-to-next-to-leading (NNLO) orders in perturbative quantum chromodynamics, and the strong coupling evolution at LO or NLO. The details about these specific MC tunes are described below:
  • 1.
    PYTHIA CP1: Tune CP1 uses the NNPDF3.1 PDF set at LO, with $\alpha_{s}$ values used for the simulation of MPI, hard scattering, FSR, and ISR equal to, respectively, 0.13, 0.13, 0.1365, and 0.1365, and running coupling according to an LO evolution CMS:2019csb .
  • 2.
    PYTHIA CP5: Tune CP5 uses the NNPDF3.1 PDF set at NNLO, with $\alpha_{s}$ values used for the simulation of MPI, hard scattering, FSR, and ISR contributions equal to 0.118, and running coupling according to an NLO evolution and the ISR emissions ordered according to rapidity CMS:2019csb .
  • 3.
    PYTHIA A3: Tune A3 uses the NNPDF2.3 PDF set at LO and running coupling according to an LO evolution QCD + QED processes. A3 tune is aimed at modelling of low-$p_{\rm T}$ QCD processes ATLAS:2016puo .
  • 4.
    PYTHIA A14: Tune A14 (ATLAS 2014) uses the NNPDF3.1 PDF set at LO and running coupling according to an LO evolution of QCD processes. This tune consists of four LO PDFs, such as CTEQ6L1, MSTW2008LO, NNPDF23LO, and HERAPDF15LO TheATLAScollaboration:2014rfk .
  • 5.
    PYTHIA CUETP8M1: A new PYTHIA 8 tune read as “CMS UE Tune PYTHIA 8 Monash set 1”. This tune uses several PDF sets and UE tunes, a more detailed discussion can be found in Ref. CMS:2015wcf
    These tunes are extracted by varying the various parameters and fitting UE observables at various collision energies CMS:2019csb , ATLAS:2016puo , TheATLAScollaboration:2014rfk .
9.2 A multi-phase transport model (AMPT)
In the present study, we use the string melting (SM) mode of AMPT (version 2.26t9b) as the particle anisotropic flow coefficients and spectra are well described at the intermediate-$p_{\rm T}$ region by the quark coalescence mechanism for hadronization Fries:2003vb , Fries:2003kq .
@@ 9.2 A multi-phase transport model (AMPT)
    Initial conditions: The AMPT model employs the heavy-ion jet interaction generator (HIJING)Wang:1991hta for initial conditions. In the HIJING model, the differential scattering cross section of the energetic minijet partons and soft strings with excited strings, which are further converted to partons, is calculated in $pp$ collisions and then parametrized into the heavy-ion collisions. These are incorporated via the nuclear overlap and shadowing function using an in-built Glauber model.
@@
    Parton transport: The produced partons are then transported into the parton transport part called Zhang’s Parton Cascade (ZPC) model Zhang:1997ej . In the String Melting version of AMPT (AMPT-SM), colored strings melt into low-momentum partons.
@@
    Hadronization: A spatial coalescence mechanism Lin:2001zk , He:2017tla is used for the hadronization of the transported partons in AMPT-SM, whereas in the default AMPT version, a Lund string fragmentation mechanism is used for hadronization of the transported partons.
@@
    Hadron transport: The final evolution of produced hadrons is done using a relativistic transport mechanism through meson-meson, meson-baryon, and baryon-baryon interactions Li:2001xh , Greco:2003mm .